# Prediction of the Most Fire-Sensitive Point in Building Structures with Differentiable Agents for Thermal Simulators

### v1

@@ Abstract
Fire safety is a critical area of research in civil and mechanical engineering, particularly in ensuring the structural stability of buildings during fire events. The Most Fire-Sensitive Point (MFSP) in a structure is the location where a fire would cause the greatest impact on structural stability. Accurate prediction of the MFSP is vital for streamlining structural assessments and optimizing the design process. This paper presents a novel framework for MFSP prediction using a neural network-based approach that integrates fire dynamics and finite element analysis through a differentiable agent model. The framework focuses on predicting the Maximum Interstory Drift Ratio (MIDR), a key indicator of structural performance under fire conditions. By leveraging the differentiable agent model, we efficiently generate labeled data for MFSP and directly train a predictor for this critical metric. To achieve this, we generated extensive simulation data encompassing structural and fire scenarios and employed graph neural networks to represent the building structures. Transfer learning was applied to optimize the training process, and an edge update mechanism was introduced to dynamically adjust edge attributes, reflecting property changes under fire conditions. The proposed model was rigorously evaluated on simulation data, demonstrating strong performance in accurately predicting both MIDR and MFSP, thus advancing fire safety analysis for building structures.
@@
Differentiable Agents, Finite Element Analysis, Fire Dynamics, Graph Neural Networks, Most Fire-Sensitive Point, Structural stability, Thermal Simulation.
footnotetext: Professor Khalid M. Mosalam is the corresponding author. This paper is currently under review at Computer-Aided Civil and Infrastructure Engineering.
@@ I Introduction
A standard fire safety requirement mandates that a building must withstand collapse for at least one hour under fire conditions at any specific point within the structure. Current methods to ensure this involve computational simulations, where fire scenarios are simulated at all potential locations within the building structure. This exhaustive approach is computationally expensive and time-consuming, increasing the cost and duration of structural design.
To address this inefficiency, we propose a novel method neural network-based framework to predict the Most Fire-Sensitive Point (MFSP) –the location where fire has the greatest impact on a structure’s stability. By identifying the MFSP, designers need to simulate only a single fire scenario to evaluate structural compliance, significantly reducing computational costs and enhancing design efficiency.
Training an MFSP predictor using traditional supervised learning is impractical due to the lack of labeled data. Generating such data requires simulating fire dynamics at every possible location within a structure –a process that is both labor-intensive and computationally prohibitive. To overcome this, we introduce an innovative framework comprising two neural network predictors: the Maximum Interstory Drift Ratio (MIDR) Predictor and the MFSP Predictor.
The workflow begins with Finite Element Analysis (FEA) simulations for a limited set of fire scenarios, which are used to train the MIDR Predictor. This predictor serves as a highly efficient and differentiable agent, simulating structural responses under fire conditions. The differentiable agent is then used in two ways: (1) as an indicator to directly train the MFSP predictor to be an “argmaxer” of the agent, and (2) to generate synthetic data to assist in the training, enabling it to efficiently identify the MFSP without extensive real-world data.
The primary contributions of this research are as follows:
  • Data Creation: We generated a large dataset of 33,000 building structures, encompassing their geometry, material properties, and nodal displacements under gravity loads. A subset of 1,573 structures was further subjected to fire scenario simulations, including nodal displacements obtained using OpenSees [21].
  • Prediction Framework: We propose a novel framework combining a differentiable agent for fire dynamics and FEA simulation with an MFSP predictor to identify the most fire-sensitive point.
@@
  • Graph Neural Networks (GNNs): We employ GNNs to model building structures and utilize Transfer Learning (TL) for efficient training. An Edge Update (EU) mechanism is introduced to account for changes in structural properties under fire scenarios.
@@
The reminder of this paper is organized as follows: Section II reviews relevant literature on fire dynamics, building structure interactions, and fire simulation tools. Section III provides an overview of the proposed system architecture. Sections IV, V, and VI detail dataset generation, the MIDR predictor, and the MSFP predictor, respectively. Implementation details and evaluation results are presented in Section VII. Finally, concluding remarks including future research directions are discussed in Section VIII. Through this research, we aim to enhance fire risk assessment accuracy and efficiency, providing a robust tool to support the design of safer, more resilient building structures.
@@ II Literature Review
Understanding the interaction between fire dynamics and building structures is fundamental to designing safer, more resilient systems. Resilience, in this context, encompasses robustness, adaptability, recoverability, and redundancy, ensuring systems can withstand, adapt to, and recover from adverse events. Recent advances in FEA, Machine Learning (ML), and physics-informed simulations have provided powerful tools for analyzing and predicting structural performance under fire conditions. This section reviews key contributions in these domains, identifies gaps, and motivates the integration of differentiable agents with thermal simulators for enhanced fire analysis in buildings.
II-A Fire Dynamics in Buildings
@@
The interplay between thermal loads and structural integrity during fire events has been extensively studied. Early work in [1] introduced a method for simulating heat transfer in steel structures. This approach was later enhanced to incorporate transient thermal responses and load redistribution effects [17], which remain vital for predicting collapse mechanisms under high-temperature conditions [10]. Despite these advances, traditional simulations often face challenges with scalability and computational efficiency in large-scale structures.
II-B Advancements in Thermal Simulation
Thermal simulation tools have evolved significantly, with recent advancements focusing on computational efficiency and accuracy. Differentiable simulators, in particular, represent a transformative paradigm by enabling gradient-based optimization for thermal and structural problems [2]. These simulators are increasingly applied in various fields, including the modeling of thermal conduction in complex structures and the optimization of structural materials for thermal resistance under uncertain environmental loads. For example, simulation tool like the Fire Dynamic Simulator (FDS) have been extensively utilized in fire engineering to model thermal and structural interactions with high fidelity [19]. Furthermore, emerging differentiable simulation frameworks, such as Taichi, provide a robust foundation for high-performance computations, enabling efficient optimization in physical modeling and material design [13]. These frameworks represent a critical leap in bridging the gap between simulation accuracy and computational efficiency.
II-C ML in Structural Analysis
Recent advancements in ML have significantly impacted structural engineering, particularly in predicting building responses under fire scenarios. For example, [23] employed modular Long Short-Term Memory (LSTM) networks to accurately predict building displacements during fire events, showcasing the potential of sequential models in time-dependent structural analysis. Similarly, GNNs, known for their strong ability to model spatially structured data [16], have gained traction in civil engineering [15]. Notably, [5] applied GNNs to predict structural responses under gravity loads, further demonstrating their efficacy in representing complex structural relationships. These developments highlight the growing role of ML techniques in enhancing the efficiency and accuracy of structural response predictions, particularly under challenging conditions like fire and other extreme events.
II-D Critical Gaps and Opportunities
Despite significant advancements, key challenges persist in structural analysis. Traditional FEA methods remain computationally intensive, limiting their scalability for large-scale simulations. Additionally, the application of GNNs in structural analysis is still in its early stages. While GNNs have been successfully employed for structural response prediction, their application to fire scenarios remains unexplored. Furthermore, GNNs have yet to fully harness global structural knowledge to identify the most vulnerable components within a structure.
A promising solution lies in treating the neural network models as differentiable agents and integrating them with thermal simulators. This approach combines the computational efficiency of ML with the physical accuracy of simulation-based methods. Expanding the use of GNNs to predict fire-sensitive points in structures could significantly improve early failure detection. Such advancements have the potential to revolutionize fire safety engineering and structural design, enhancing both resilience and efficiency.
@@ III System Overview
The primary goal of this work is to accurately predict the MFSP in a building structure subjected to various fire scenarios. To achieve this, we propose an integrated framework that combines GNNs and FEA. The system architecture is illustrated in Figure 1 and comprises two key components: the Maximum Interstory Drift Ratio (MIDR) Predictor and the MFSP Predictor. This framework seamlessly integrates physics-based simulations, GNN-driven predictive modeling, and data-driven techniques, enabling efficient and accurate MFSP prediction, providing a powerful tool for proactive fire safety analysis and risk mitigation in building structures.
Figure 1: Overview of the proposed framework for predicting the MFSP in building structures.
III-A MIDR Predictor
@@
A GNN processes this input to represent the structural configuration and fire location as a graph. The MIDR Predictor is trained on labeled data generated using OpenSees, a robust open-source FEA framework [21]. These labels represent detailed structural responses under various fire conditions. Once trained, the MIDR Predictor functions as a differentiable agent, offering computationally efficient MIDR estimates. Its capabilities include:
@@
  • Integrating with Neural Networks: Reducing the computational cost typically associated with simulation-based methods.
III-B MFSP Predictor
The MFSP Predictor acts as an “argmaxer module” for the MIDR Predictor, identifing the fire location that results in the highest MIDR. This location, deemed the MFSP, corresponds to the point of greatest structural vulnerability. By leveraging the structural graph as input and utilizing the MIDR Predictor’s outputs, the MFSP Predictor efficiently pinpoints the critical fire location.

### v5

@@ Abstract
Fire safety is crucial for ensuring the stability of building structures, yet evaluating whether a structure meets fire safety requirement is challenging. Fires can originate at any point within a structure, and simulating every potential fire scenario is both expensive and time-consuming. To address this challenge, we propose the concept of the Most Fire-Sensitive Point (MFSP) and an efficient machine learning framework for its identification. The MFSP is defined as the location at which a fire, if initiated, would cause the most severe detrimental impact on the building’s stability, effectively representing the worst-case fire scenario. In our framework, a Graph Neural Network (GNN) serves as an efficient and differentiable agent for conventional Finite Element Analysis (FEA) simulators by predicting the Maximum Interstory Drift Ratio (MIDR) under fire, which then guides the training and evaluation of the MFSP predictor. Additionally, we enhance our framework with a novel edge update mechanism and a transfer learning-based training scheme. Evaluations on a large-scale simulation dataset demonstrate the good performance of the proposed framework in identifying the MFSP, offering a transformative tool for optimizing fire safety assessments in structural design. All developed datasets and codes are open-sourced online.
@@
Differentiable agents, Finite Element Analysis, Graph Neural Networks, Most Fire-Sensitive Point, Structural stability, Thermal simulation.
footnotetext: Professor Khalid M. Mosalam is the corresponding author. This paper has been accepted by and is to appear in Computer-Aided Civil and Infrastructure Engineering.
@@ I Introduction
In designing a building structure, there are many requirements to be met, e.g., requirements related to gravity load stability [2], earthquake resistance, wind effects, and fire safety [14]. Before a final check with multiple load combinations, e.g., combination of gravity, earthquake and fire loads, each individual requirement should be satisfied in a pre-check phase. A standard fire safety requirement mandates that a building must withstand collapse for at least one hour under fire conditions at any specific point within the structure. Current methods to ensure this requirement involve computational simulations, e.g., using Finite Element Analysis (FEA) tools, where fire scenarios are simulated at all potential locations within the building structure. This exhaustive approach is computationally expensive and time-consuming, greatly hindering the iterative process in the preliminary design of structures.
To address this inefficiency, we first propose the concept of the Most Fire-Sensitive Point (MFSP) –the location where fire has the greatest impact on a structure’s lateral stability. The MFSP corresponds to the worst-case scenario when a fire occurs. By identifying the MFSP, designers may only need to simulate its corresponding fire scenario to evaluate the structural compliance in the pre-check phase, significantly reducing computational costs and effectively streamlining the design process. However, identifying the MFSP is not a trivial task –conventionally, brute-force methods with computational simulations are still required. Fortunately, recent Machine Learning (ML) advancements provide new opportunities to efficiently identify the MFSP as conducted in this study. Regarding the fire scenarios, [31] and [43] employed Recurrent Neural Network (RNN) for real-time displacement prediction, focusing on post-fire collapse warning systems by forecasting future outcomes through early-stage fire features. On the other hand, [6] utilized Graph Neural Network (GNN) for post-fire concrete temperature field mapping with material-level assessment. Despite these works, using ML techniques to assess the overall structural lateral stability in the design stage without real-time information of fires remains a major challenge.
Moreover, even with a Neural Network (NN) predictor, it is difficult to directly apply traditional supervised learning for MFSP prediction. Traditional supervised learning requires large amount of labeled data, i.e., structures with ground truth MFSP labels. However, obtaining such labeled data is impractical due to the required brute-force simulations, which are both labor-intensive and computationally prohibitive. To overcome this challenge, we propose a framework with two GNN-based predictors: the Maximum Interstory Drift Ratio (MIDR) predictor as the differentiable agent of fire simulations to assess the overall structural lateral stability, and the MFSP predictor as the “argmaxer” of the MIDR predictor to predict the MFSP.
The workflow begins with FEA simulations for a limited set of fire scenarios, which are used to train the MIDR predictor. In traditional FEA simulations, the complex operations due to nonlinear material behavior, iterative solvers, and discrete time stepping make traditional FEA tools non-differentiable as well as computationally expensive. In contrast, our differentiable agent uses a GNN with differentiable operations (e.g., a ReLU-activated Multi-Layer Perceptron (MLP)), ensuring computations of the gradients for end-to-end training. Compared to traditional physics-based simulations, the GNN-based MIDR predictor offers two major advantages: (1) the differentiability enables the agent to work as an indicator to directly guide the training of the MFSP predictor to be an “argmaxer” of the agent, and (2) the computational efficiency enables it to generate large synthetic data to assist in the training, allowing it to efficiently identify the MFSP without requiring extensive real-world data.
This study focuses on framed steel building structures, which are widely used in modern construction, including commercial facilities. However, the proposed research method is not limited to steel structures and can be extended to other types of buildings with similar structural configurations. The primary contributions of this research are as follows:
  • Concept of MFSP & workflow for its prediction: We first propose the concept of MFSP, which can greatly enhance the structural design process by reducing time and computational cost to check whether fire requirements are satisfied. We also propose a novel workflow to predict the MFSP for a building structure, including a differentiable neural agent for fire dynamics and FEA simulation and an MFSP predictor trained with the aid of the agent.
@@
  • Transfer Learning (TL) & Edge Update (EU) in GNNs: We propose a GNN-based framework to model building structures and incorporate TL for efficient training. An Edge Update (EU) mechanism is introduced to account for the changes in the material properties under fire scenarios.
  • Dataset generation & verification of proposed workflow: We generate a large-scale dataset of building structures, encompassing their geometry, material properties, and nodal displacements under gravity loads. After filtering out unreasonable structures in the whole dataset, 16,050 structures form the unlabeled dataset, and another 1,573 structures are simulated for 30 fire scenarios, forming the labeled dataset, which contains the nodal displacements for each of the $1,573\times 30=47,190$ fire cases, obtained using a FEA tool, namely, $\chi$ara [33] (previously named OpenSeesRT [32]). Simulation results based on the generated dataset verify the efficiency and accuracy of the proposed workflow. Ablation experiment also demonstrates the effectiveness of the proposed TL and EU mechanisms.
@@
After this introduction, the paper is organized as follows: Section II reviews relevant literature on GNN applications, agents, and fire analysis of buildings. Section III provides an overview of the proposed system architecture. Sections IV, V, and VI detail the MIDR predictor, the MSFP predictor, and the dataset generation, respectively. Implementation details and evaluation results are presented in Section VII. We discuss the limitations and proposed future work in Section VIII. Finally, Section IX is a concise summary of the main conclusions of the study. Through this research, we aim to enhance fire risk assessment accuracy and efficiency, providing a robust tool to support the design of safer, more resilient buildings.
@@ II Literature Review
Understanding the interaction between fire dynamics and building structures is fundamental to designing safer, more resilient building systems. Resilience, in this context, encompasses robustness, adaptability, recoverability, and redundancy, ensuring systems can withstand, adapt to, and recover from adverse events. Recent advances in FEA, ML, and physics-informed simulations have provided powerful tools for analyzing and predicting the structural performance under fire conditions. This section reviews key contributions in these domains, identifies gaps, and motivates the integration of differentiable agents with thermal simulators for enhanced fire analysis in buildings.
II-A GNN Applications and Agents
Recently, GNN has shown successful applications in many areas in civil engineering. For example, [20] employed GNN for predicting network-wide metro passenger flow with dynamic propagation graphs, and [37] introduced multi-graph convolutional network designed to predict the future trajectories of traffic agents near an autonomous vehicle by incorporating various interaction perspectives and EGO-planning information, where EGO stands for Euclidean Signed Distance Field-free Gradient-based lOcal. Moreover, [15] utilized physics-informed GNN-assisted auto-encoder to reconstruct high-resolution urban wind fields based on sparse sensor data. [23] reviewed the application of GNNs in construction. These successful applications demonstrate the strong ability of GNN to capture spatial relationships and the potential for applying GNN in structural analysis. In predicting structural responses, [8] applied GNN to a static problem to predict responses such as nodal displacements and shear forces, showing very high accuracy. GNN was further combined with RNN in [22] to predict building thermal load in general scenarios, focusing on capturing spatial interactions between multiple thermal zones and temporal dependencies.
Despite these successful applications, using GNN to predict the structural response in fire scenarios has not been well studied. [43] proposed a GNN-RNN hybrid model for real-time displacement prediction of steel frames under fire, yet the work focused on post-fire collapse warning systems by forecasting future outcomes through early-stage fire features, rather than evaluating the stability at the design phase. Meanwhile, [6] developed a GNN-based model for post-fire concrete temperature field mapping, but their approach remained limited to material-level assessment and did not address the response at the structural level. Moreover, existing works merely use GNN for inference, i.e., prediction, neglecting that it could also work as an agent for other purposes. The concept of agent has been widely used in civil engineering applications. For example [16] proposed a multi-agent replicator controller for vibration control of smart structures, and [44] introduced a multi-agent Reinforcement Learning (RL) model for optimizing maintenance decisions in interdependent highway pavement networks. However, these works mainly treat agents as the final objective, i.e., the well-trained agents are directly used to generate the desired decisions. We distinguish the presented study from these works by the role of the agent in the framework, where herein the agent works as a surrogate model and is an intermediate module in the proposed framework. The agent’s differentiability and high efficiency are fully utilized in the presented work.
II-B Thermal Simulation
@@
Fire event simulation encompasses several critical aspects, including heat transfer modeling, fire spread dynamics, and the impact of thermal loads on structural elements. The interplay between thermal loads and structural elements during fire events has been extensively studied. Early work in [3] introduced a method for simulating heat transfer in steel structures. This approach was later enhanced to incorporate transient thermal responses and load redistribution effects [25], which remain vital for predicting collapse mechanisms under high-temperature conditions [13].
Beyond heat transfer analysis, fire simulation also involves assessing the deterioration of structural resistance. This includes temperature-induced material degradation, progressive weakening of load-bearing components, and failure mechanisms under prolonged fire exposure. Notably, steel structures experience reductions in Young’s modulus and yield strength at elevated temperatures, which affect their ability to sustain applied loads. Thermal expansion and differential heating further contribute to internal stress redistribution, potentially leading to local or global instability. Although phenomena such as buckling of slender members, loss of composite action in beam-column connections, and creep effects under sustained high temperatures are integral considerations in advanced fire analysis models, they are beyond the scope of the present paper and should be considered in future studies.
Recently, [42] developed a computational framework to enhance the simulation of welding residual stresses using ML and optimization techniques, validated through experimental measurements. [41] studied the thermal contraction coordination behavior between an unbound aggregate layer and an asphalt mixture overlay using a finite difference and discrete element coupling method. Most past works focused on local behavior of system components under thermal loads instead of the overall structural stability.
Computational Fluid Dynamics (CFD)-based methods, such as Fire Dynamics Simulator (FDS) [28], provide high-fidelity modeling of temperature evolution, flame spreading, and gas-phase combustion. However, the computational cost remains a significant limitation, restricting large-scale structural assessments. FDS requires not only a large computing time for the simulation, but also significant modeling time for the building structures. This limits the use of FDS to case studies with necessary input of building structural details, as well as information on combustible materials inside the building. For example, [27] employed FDS modeling to investigate two ignition scenarios in 2014 Romanian restaurant fire, determining the most likely cause of rapid fire spread and lethal Carbon Monoxide exposure that resulted in four fatalities.
Real-world fire loads often display substantial spatial variability across different rooms in a building [10], resulting in scenario-specific temperature fields with limited generalizability. Furthermore, even within identical scenarios, variations in fire modeling methodologies can produce distinctly different temperature fields, as demonstrated by [46] using a full-scale experiment to develop a new temperature field prediction model for large space fires, and by [9] through creating a new temperature-time curve to better represent the non-uniform temperature distribution in large space fires compared to standard curves. Moreover, studies on bridge fires [18] demonstrated that environmental factors, such as wind speeds, can significantly influence the temperature distributions. These challenges emphasize the need for efficient and adaptive methods to generate fire temperature data, where less important factors might be neglected.
While existing studies have extensively explored post-fire structural behavior, integrating material deterioration mechanisms into a ML-based predictive framework remains an open challenge. The proposed approach in this study aims to bridge this gap by efficiently approximating fire-induced structural responses without relying on computationally intensive physics-based simulations. To overcome the above difficulties, we propose a rule-based temperature generation method, which is efficient and scalable for serving large-scale dataset simulations. With a generated temperature field in a structure, FEA tools, e.g., [45], can be used to obtain the structural response under thermal loads. For steel structures, previous studies, e.g., [31], have demonstrated that they rapidly equilibrate with the surrounding gas temperatures due to efficient heat exchange. Consequently, gas temperatures can be directly used as input for FEA. This inspired us to generate a large dataset using FEA for thermal analysis in fire scenarios to form the foundation of the NN training and testing.
@@ III System Overview
The primary goal of this work is to accurately predict the MFSP in building structures subjected to various fire scenarios by utilizing the MIDR as a metric for the overall lateral stability. To achieve this, we propose an integrated framework that combines GNNs and FEA. The system architecture is illustrated in Figure 1 and comprises two key components: the MIDR predictor and the MFSP predictor. There are two stages corresponding to the MIDR predictor’s different modes. In stage 1, we use data with ground truth MIDR values obtained from FEA simulations to train the MIDR predictor. Then, in stage 2, with the well-trained MIDR predictor working as a differentiable agent of fire simulation, we train the MFSP predictor to determine the point that maximizes the MIDR. This framework seamlessly integrates physics-based simulations, GNN-driven predictive modeling, and data-driven techniques, enabling efficient and accurate MFSP prediction to provide a powerful tool for proactive fire safety analysis and risk mitigation in building structures.
Figure 1: Proposed framework for predicting the MFSP in building structures. Trapezoids and gray rectangles represent the NN predictor and processing, respectively. Dashed and solid lines indicate that MIDR predictor is in the respective training mode to determine its parameters and evaluation mode with parameters fixed, acting as a differentiable agent.
III-A Structural Stability Metrics
The Interstory Drift Ratio (IDR) of each node serves as a critical parameter for evaluating structural lateral stability and deformation under external forces. IDR quantifies the relative displacement between two consecutive floors (interstory displacement) as a percentage of the floor height. Mathematically, the IDR for a given node $i$ is defined as follows:
where the numerator represents the relative displacement (in the horizontal plane $X-Y$) of node $i$ with respect to the corresponding node on the floor below, and $H$ denotes the story height between these two floors. Excessively high drift ratios indicate significant structural deformation, potentially leading to major damage or collapse due to lateral instability. The Maximum IDR among all the nodes of a structure, i.e., MIDR, is chosen to be a representative example metric for assessing the overall structural stability performance during fire events. Although MIDR may not capture all possible failure modes, such as local collapses due to midspan softening, local buckling, or loss of vertical elements, we emphasize that the proposed method is not limited to MIDR. The framework can be adapted to other performance indicators of interest, by simply replacing the MIDR with the desired metric.
Remark: Selecting MIDR as the primary metric for assessing the structural integrity under fire conditions is motivated by its effectiveness in quantifying global deformation patterns. In fire-induced scenarios, thermal expansion, stiffness degradation, and gravity-induced deformations contribute to structural instability, which can be captured through relative floor displacements. MIDR provides a direct and interpretable measure of structural vulnerability by identifying floors experiencing excessive lateral deformations that may lead to global instability or loss of vertical load-carrying capacity. While MIDR is commonly used for seismic and wind-induced responses, its application to fire scenarios is justified as a fire-driven thermal effect also leads to large-scale deformation patterns, particularly in multi-story steel structures. Importantly, MIDR serves as an effective proxy for overall building stability in computational frameworks where parameterizing every potential failure mode (e.g., local buckling, connection failure, progressive collapse) is not feasible. However, fire-induced failure mechanisms extend beyond interstory drift, and localized effects are not explicitly captured by MIDR. These mechanisms typically develop locally and may not always translate into immediate global structural instability. While our current framework focuses on identifying the MFSP based on a worst-case drift metric, future work could integrate alternative failure criteria to further refine the fire vulnerability predictions.
Note that “M” in MIDR and MFSP represents different concepts. In the case of MIDR, for a given structure and fire source point, the IDR is computed at each node, and the MIDR is defined as the maximum IDR among all nodes. In contrast, MFSP refers to the fire source location that results in the highest MIDR across all possible fire source points within the structure. In summary, an MIDR is associated with a specific structure and fire source point pair, whereas an MFSP characterizes an entire structure by identifying the most critical fire source location.
III-B MIDR & MFSP Predictors
@@
A GNN processes this input to represent the structural configuration and fire location as a graph. The MIDR predictor is trained on labeled data generated using $\chi$ara [33], a robust open-source FEA framework based on OpenSees [24]. These labels represent detailed structural responses under various fire conditions. Once trained, the MIDR predictor functions as a differentiable agent, offering computationally efficient MIDR estimates. Its capabilities include:
@@
  • Integrating with NNs: Reducing the computational cost typically associated with simulation-based methods.
The MFSP predictor acts as an “argmaxer module” for the MIDR predictor, identifing the fire location that results in the highest MIDR. This location corresponds to the point of the greatest structural vulnerability. By leveraging the structural graph as input and utilizing the MIDR predictor’s outputs, the MFSP predictor efficiently pinpoints the critical fire location.